# PyTomography: A Python Library for Medical Image Reconstruction

### v4

@@ 3 Illustrative examples
Multiple examples are explored to highlight the capabilities of PyTomography. Specifically, reconstruction of parallel-collimator SPECT and LM-TOF PET data is demonstrated, but it should be emphasized that these are only intended as example use cases for particular imaging modalities. Validation against other software libraries is also shown in the Appendix, which contains (i) SIMIND simulated SPECT data reconstruction validated against STIR, (ii) DICOM SPECT data (obtained on a GE Tandem Discovery 670 Pro) validated against MIM’s SPECT reconstruction functionality, and (iii) GATE PET data reconstruction validated against CASToR. We chose to validate DICOM reconstruction against MIM, and not the Siemens internal reconstruction software, because it represents an external library with access to the same information as PyTomography.
@@ 3.1 SPECT
This section highlights SPECT reconstructions using available reconstruction algorithms in PyTomography, demonstrating differences in quantitative metrics between different algorithms. Two examples are considered: (i) reconstruction of Monte Carlo SPECT data generated via SIMIND, and (ii) reconstruction of clinical SPECT data. In the clinical case, the patient received radiopharmaceutical therapy with 177Lu-PSMA for prostate cancer and was imaged 3 hours post injection; the XCAT phantom used as input for SIMIND was configured such that the kidney, liver, and lesion activity concentrations roughly approximated the clinical patient. In each case, the data correspond to acquisition on a GE Tandem Discovery 670 Pro SPECT system (MEGP collimators) with 60 projections (25 s per projection). The clinical data consists of acquisition at two bed positions with projections of shape $128\times 128$, while the simulated data consists of acquisition at one bed position with projections of shape $128\times 240$. The two lesions inserted in the XCAT phantom were (i) based on an observed bone metastasis in the sternum of the clinical patient, and (ii) placed in the liver. In each case, the system matrix employed attenuation and PSF modeling, and the likelihood function considered an additive scatter term obtained via the triple energy window technique.
The reconstruction algorithms considered were OSEM (4it/6ss), BSREM-RDP (40it/6ss), BSREM-RDPAP (40it/6ss), and KEM (10it/6ss) [’it’ for iterations; and ’ss’ for subsets]; algorithms using RDP used the hyperparameters $\beta=0.3$, $\gamma=2$, RDPAP only considered the 8 nearest neighbours based on the attenuation map, and KEM used Eq. 5 of Vuohijoki [6] for its filter. The time required for reconstruction of one bed position of the clinical data was 1.07 s/it (OSEM), 0.83 s/it (BSREM-RDP), 0.86 s/it (BSREM-RDPAP), and 1.07 s/it (KEM). The time required for reconstruction of the simulated data was 1.32 s/it (OSEM), 1.13 s/it (BSREM-RDP), 1.21 s/it (BSREM-RDPAP), and 1.52 s/it (KEM).
@@
For the clinical data, the left kidney was segmented using the CT-based TotalSegmentator model [24], and the bone metastasis in the sternum was segmented by first segmenting the lesion from the OSEM reconstruction, and then applying a threshhold mask greater than 0.150.150.15~{}0.15cm-1 so that only voxels within the bone were considered. A spherical mask was used in the liver for both the simulated and clinical data for measurement of noise in each reconstruction algorithm; noise was quantified as a percent by computing the standard deviation of uptake divided by the mean of uptake in that region. Reconstructions of the clinical and simulated data are shown in Figures 2 and 3 respectively.
@@
Figure 2: Top: Maximum intensity projections of reconstructed clinical SPECT data; the leftmost subplot corresponds to reconstruction using the MIM reconstruction software library. Three annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL. Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
Figure 3: Top: Maximum intensity projections of reconstructed simulated SPECT data; the leftmost subplot corresponds to the ground truth XCAT phantom used for simulation in SIMIND. Four annotations are shown: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, and (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL, and (iv) the location of the liver metastasis (purple arrow) with the corresponding activity concentration in kBq/mL. Bottom: Axial slice corresponding the hottest slice of the bone metastasis; SPECT uptake is shown in color and the attenuation map is shown in greyscale.
@@ 3.2 PET
    Use of the DIPRecon algorithm to enhance reconstruction
The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps; this fictional scanner is representative of a Siemens Biograph mMR with TOF capabilities. The digital phantom used for simulation in GATE consisted of an ultra-high resolution 18F-FDG PET image, a T1-weighted MRI brain image, and a corresponding attenuation map [25]. The total uptake of 15.66 MBq was representative of a typical 18F-FDG distribution 1 hour post injection. The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries. In all cases, simulated data were reconstructed using a $128\times 128\times 96$ matrix size with voxel spacing of 222~{}2mm; 5 TOF bins were used with a total range of 30 cm. Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction. The “scatters” contribution for each listmode event was estimated using the TOF single scatter simulation (SSS) capabilities of the library, based on Watson [26]. The PET image used as a proxy for SSS was a reconstruction of the data via LM-TOF OSEM (40it, 1ss) without scatter estimation. Simulation and reconstruction were then performed in the following three ways:
@@
    A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan. The data was then reconstructed using LM-TOF OSEM (80it, 1ss) with attenuation, random, and scatter modeling. PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 4 mm.
@@
    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed.
@@
    A low count reconstruction was performed without PSF modeling and used in the DIPRecon algorithm; no PSF modeling was used because (i) the DIP performs better on images with high frequency noise, and (ii) the DIPRecon algorithm already has noise reduction properties. The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the low count reconstruction. The neural network was a UNet [27] with similar architecture to the original DIP paper [8]; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code.
The time required for the complete TOF-SSS estimation was 665 s (high count) and 577.5 s (low count), the time required for the OSEM reconstructions were 3.31 s/it (high count) and 1.18 s/it (low count), and the time required for the deep image prior reconstruction was 289.1 s (initial fit) and 21.53 s/it (reconstruction). Corresponding phantom reconstructions, as well as bias and noise in grey and white matter as a function of iteration number, are shown in Figure 4.
Figure 4: Left: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [25]. From top to bottom: axial, coronal, and sagittal slices. From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (80it/1ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction. Right: Variation in absolute bias and standard deviation throughout iterative reconstruction. Subplots correspond to the grey matter region (upper) and white matter region (lower). For the low count rate scenario, points are plotted every 10 iterations. For the low count rate scenario with OSEM, the first point corresponds to the 10th iteration. The bias was initially positive and became negative for all three cases in white matter.
@@ 4 Impact
In this work, we have motivated and presented the software architecture of PyTomography, and showed use cases on both SPECT and PET data. The short times required for reconstruction using this tool can enable researchers to perform extensive studies on digital/physical phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms. To demonstrate (i) the impact of the software and (ii) the extent to which new research questions can be pursued, we will discuss each of the illustrative examples.
A variety of algorithms were examined for 177Lu SPECT reconstruction; various quantitative metrics were estimated to highlight the differences between the available algorithms in the library. For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP. Furthermore, the simulated data prediction from BSREM-RDPAP (807.02.17807.02.17807.02.17~{}807.02.17kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration. For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise. BSREM-RDPAP has a lower noise (6.9%) than OSEM (13.3%) but with higher uptake in the kidney (807.02 kBq/mL for BSREM-RDPAP vs. 755.07 kBq/mL for OSEM) for the simulated data; BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction. For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown on Figures 2 and 3. These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (217.13217.13217.13~{}217.13kBq/mL for BSREM-RDPAP and 216.15216.15216.15~{}216.15kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (198.88198.88198.88~{}198.88kBq/mL). For the liver lesion in the simulated data, however, use of the anatomical reconstruction algorithms decreased the activity concentration predictions (194.04194.04194.04~{}194.04kBq/mL for BSREM-RDPAP and 231.31231.31231.31~{}231.31kBq/mL for KEM) compared to OSEM (241.08241.08241.08~{}241.08kBq/mL). This occured because there is less anatomical difference in the surrounding tissue in the liver metastasis compared to the bone metastasis, so the smoothness prior merely blurred and reduced the activity concentration. These results, which are consistent with previous studies [6], suggest that anatomical priors may yield more accurate activity concentrations for bone lesion dosimetry. While this example focused primarily on organ and lesion dosimetry, PyTomography users will be to assess the merits of various reconstruction algorithms on any task specific metric.
@@
While this paper has demonstrated implementations for parallel collimator SPECT and LM-TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, various forms of computed tomography, (CT), magnetic resonance (MR) imaging, and compton camera (CC) imaging. As displayed in Figure 1, projectors for different imaging modalities can be implemented using a variety of python based tools: it is hoped that the structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems. Implemented code does not necessarily need to be written directly in python; the parallelproj library used for PET modeling, for example, uses a CUDA C backend for its calculations. This flexibility permits contributions from users familiar with a variety of coding languages. Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own data. The immediate goals for future development in PyTomography at the time of publication include:
@@
    Fast PSF modeling for high energy SPECT isotopes, such as 131I and 225Ac.
@@
    Expanded support for clinical PET data.
@@
In summary, we hope that we have made a compelling case for the benefit that this tool can bring to the medical imaging community. It can be used as a framework for studies that require fast reconstruction times, such as those that explore a wide range of reconstruction hyperparameters. It may also be used by developers seeking to implement new system matrices and likelihoods for their own imaging systems. We highly encourage and welcome open-source contributions to assist in the items listed above; those who wish to contribute are encouraged to read the developers’ guide on the documentation website.

### v5

@@ 3 Illustrative examples
Multiple examples are explored to highlight the capabilities of PyTomography. Specifically, reconstruction of parallel-collimator SPECT and LM-TOF PET data is demonstrated, but it should be emphasized that these are only intended as example use cases for particular imaging modalities. Validation against other software libraries is also shown in the Appendix, which contains (i) SIMIND simulated SPECT data reconstruction validated against STIR, (ii) DICOM SPECT data (obtained on a GE Tandem Discovery 670 Pro) validated against MIM’s SPECT reconstruction functionality, and (iii) GATE PET data reconstruction validated against CASToR. We chose to validate DICOM reconstruction against MIM, and not the GE internal reconstruction software, because it represents an external library with access to the same information as PyTomography.
@@ 3.1 SPECT
This section highlights SPECT reconstructions using available reconstruction algorithms in PyTomography, demonstrating differences in quantitative metrics between different algorithms. Two examples are considered: (i) reconstruction of clinical SPECT data and (ii) reconstruction of Monte Carlo SPECT data generated via SIMIND. In the clinical case, the patient received radiopharmaceutical therapy with 177Lu-PSMA-617 for prostate cancer and was imaged 3 hours post injection; the XCAT phantom used as input for SIMIND was configured such that the kidney, liver, and lesion activity concentrations roughly approximated the clinical patient. In each case, the data correspond to acquisition on a GE Tandem Discovery 670 Pro SPECT system (MEGP collimators) with 60 projections (25 s per projection). The clinical data consists of acquisition at two bed positions with projections of shape $128\times 128$, while the simulated data consists of acquisition at one bed position with projections of shape $128\times 240$. The two lesions inserted in the XCAT phantom were (i) based on an observed bone metastasis in the sternum of the clinical patient, and (ii) placed in the liver. In each case, the system matrix employed attenuation and PSF modeling, and the likelihood function considered an additive scatter term obtained via the triple energy window technique. The reconstruction algorithms considered were OSEM (4it/6ss), BSREM-RDP (40it/6ss), BSREM-RDPAP (40it/6ss), and KEM (10it/6ss) [‘it’ for iterations; and ‘ss’ for subsets]; algorithms using RDP used the hyperparameters $\beta=0.3$, $\gamma=2$, RDPAP only considered the 8 nearest neighbours based on the attenuation map, and KEM used Eq. 5 of Vuohijoki [6] for its filter. The required computation times are shown in Table 3.
@@
For the clinical data, the left kidney was segmented using the CT-based TotalSegmentator model [27], and the bone metastasis in the sternum was segmented by first segmenting the lesion from the OSEM reconstruction, and then applying a threshhold mask greater than 0.150.150.15~{}0.15cm-1 so that only voxels within the bone were considered. A spherical mask was used in the liver for both the simulated and clinical data for measurement of noise in each reconstruction algorithm; noise was quantified as a percent by computing the standard deviation of uptake divided by the mean of uptake in that region. Reconstructions of the clinical and simulated data are shown in Figure 2.
@@
Figure 2: DICOM SPECT reconstruction of clinical data (top) and SIMIND SPECT reconstruction of Monte Carlo data (bottom). Greyscale images correspond to maximum intensity coronal projections of reconstructed images. The annotations correspond to: (i) the left kidney, contained inside a pink box, with the corresponding activity concentration in kBq/mL (ii) the location of the spherical mask in the liver used for noise quantification with the corresponding noise in percent, (iii) the location of the bone metastasis (blue arrow) with the corresponding activity concentration in kBq/mL, and (iv) the location of the liver lesion (purple arrow) inserted in the SIMIND data. Axial slices are shown below each greyscale image that correspond to the slice containing the bone metastasis; the SPECT activity is shown with a magma colormap superimposed on a Greyscale CT image.
@@ 3.2 PET
    Use of the DIPRecon algorithm to enhance reconstruction; the DIPRecon class of PyTomography directly takes as input an instance of an nn.Module of PyTorch, highlighting the library capabilities to interface directly with neural networks.
The simulated scanner consisted of $8$ rings of 338 mm radius with 32.25 mm ring spacing, and a TOF coincidence resolution of 550 ps; this fictional scanner is representative of a Siemens Biograph mMR with TOF capabilities. The digital phantom used for simulation in GATE consisted of an ultra-high resolution 18F-FDG PET image, a T1-weighted MRI brain image, and a corresponding attenuation map [28]. The total uptake of 15.66 MBq was representative of a typical 18F-FDG distribution 1 hour post injection. The normalization $\eta$ was obtained via simulation of a thin cylindrical shell of radius 318318318~{}318mm that extended along the axial FOV of the scanner; normalization computation is available in PyTomography for basic polygonal PET geometries. In all cases, simulated data were reconstructed using a $204\times 204\times 154$ matrix size with voxel spacing of 1.251.251.25~{}1.25mm; 5 TOF bins were used with a total range of 30 cm. Before reconstruction, the “randoms” contribution to each listmode event was estimated from delayed coincidences; this process consisted of (i) binning the delayed coincidence events into a sinogram (ii) smoothing the sinogram to reduce noise, and (iii) converting back into LM form for use in LM reconstruction. The “scatters” contribution for each listmode event was estimated using the TOF single scatter simulation (SSS) capabilities of the library, based on Watson [29]. The PET image used as a proxy for SSS was a reconstruction of the data via LM-TOF OSEM (40it, 1ss) without scatter estimation. Simulation and reconstruction were then performed in the following three ways:
@@
    A 9 minute acquisition (high count) was simulated in GATE that included all relevant physical phenomenon in a typical PET scan. The data was then reconstructed using LM-TOF OSEM (4it, 20ss) with attenuation, random, and scatter modeling. PSF modeling was implemented via 3 dimensional blurring in object space using a Gaussian kernel with a full width half max of 3 mm.
@@
    The dataset above was subsampled into a 3 minute acquisition (low count), and an identical reconstruction was performed except with 80it and 1ss.
@@
    A low count reconstruction was performed without PSF modeling and used as an initial image for the DIPRecon algorithm; no PSF modeling was used in the initial image because (i) the DIP performs better on images with high frequency noise, and (ii) the DIPRecon algorithm already has noise reduction properties. The prior image $z$ used was the corresponding T1-weighted MRI image; the initial network prediction $\mathcal{N}(z|\theta_{0})$ was obtained via fitting to the low count reconstruction. The neural network was a UNet [30] with similar architecture to the original DIP paper [8]; more information on the network structure and reconstruction hyperparameters can be found in the publicly available code. Although PSF modeling was not used to obtain the initial image, it was employed during the iterative process of the DIPRecon algorithm.
Corresponding phantom reconstructions, as well as bias and noise in grey and white matter as a function of iteration number, are shown in Figure 3. Required computation times are shown in Table 3.
Figure 3: Top: PET reconstruction of GATE simulated data corresponding to a ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [28]. From top to bottom: axial, coronal, and sagittal slices. From left to right: ground truth MRI (non contrast adjusted), ground truth PET uptake, high count OSEM reconstruction (4it/20ss), low count OSEM reconstruction (80it/1ss), and low count DIP reconstruction. Bottom: Variation in absolute bias and standard deviation throughout iterative reconstruction. Subplots correspond to the grey matter region (left) and white matter region (right). For the low count rate scenario, points are plotted every 10 iterations, and for the low count OSEM example, the first point corresponds to the 10th iteration. The bias was initially positive and became negative for all three cases in white matter.
Table 3: Computational times for all use cases considered in this paper. Times required to open raw data are not considered. Times to construct SPECT system matrices and compute PET randoms in PyTomography are not considered since these are negligible. The notation “s / it” corresponds to the seconds per iteration for a given reconstruction algorithm, ‘SSS’ refers to the single scatter simulation in PET, and ‘SM’ refers to the creation of the system matrix. For the DIPRecon PET case, ‘LC’ refers to the prior low count reconstruction, since this is required to obtain an initial image for DIPRecon.
SPECT DICOM Time Required
OSEM (4it/6ss) (1.32 s/it/bed) $\times$ 4it $\times$ 2 bed $=$ 10.6 s
BSREM-RDP (40it/6ss) (1.13 s/it/bed) $\times$ 40it $\times$ 2 bed $=$ 90.4 s
BSREM-RDPAP (40it/6ss) (1.21 s/it/bed) $\times$ 40it $\times$ 2 bed $=$ 96.8 s
KEM (10it/6ss) (1.52 s/it/bed) $\times$ 10it $\times$ 2 bed $=$ 30.4 s
SPECT SIMIND Time Required
OSEM (4it/6ss) (1.07 s/it) $\times$ 4it $=$ 4.3 s
BSREM-RDP (40it/6ss) (0.83 s/it) $\times$ 40it $=$ 33.2 s
BSREM-RDPAP (40it/6ss) (0.86 s/it) $\times$ 40it $=$ 34.4 s
KEM (10it/6ss) (1.07 s/it) $\times$ 10it $=$ 10.7 s
PET GATE Time Required
High Count OSEM (4it/20ss) $(16.9~{}\text{s/it})&\times 4\text{it}+37.4~{}\text{s (SM)}\\ +&343.8~{}\text{s (SSS)}=\textbf{449.0~{}\text{s}}\end{aligned}$
Low Count OSEM (80it/1ss) $(3.15~{}\text{s/it})&\times 80\text{it}+30.8~{}\text{s (SM)}\\ +&375.7~{}\text{s (SSS)}=\textbf{658.3~{}\text{s}}\end{aligned}$
Low Count DIPRecon (100it) $(21.5~{}\text{s/it})&\times 100\text{it}+658.3~{}\text{s (LC)}\\ =&\textbf{2808.5~{}\text{s}}\end{aligned}$
SPECT Validation Time Required
PyTomo OSEM (2it/8ss) (2.70 s/it) $\times$ 2it $=$ 5.4 s
STIR OSEM (2it/8ss) 51542.2 s
PET Validation Time Required
PyTomo OSEM (2it/21ss) (59 s/it) $\times$ 2it $=$ 118 s
CASToR OSEM (2it/21ss) (455 s/it) $\times$ 2it $=$ 910 s
@@ 4 Impact
In this work, the software architecture of PyTomography was demonstrated and use cases were explored on both SPECT and PET data. The efficient reconstruction times, shown in Table 3, can enable researchers to perform extensive studies on digital/physical phantom and patient data that may include multiple patients, noise realizations, and reconstruction algorithms. A detailed discussion on each use case is included to highlight example analyses that can be performed with the software.
A variety of algorithms were examined for 177Lu SPECT reconstruction; various quantitative metrics were estimated to highlight the differences between the available algorithms in the library. For both the simulated and clinical data, the algorithms yielding smallest to largest kidney concentration were BSREM-RDP, OSEM, KEM, and BSREM-RDPAP. The simulated data prediction from BSREM-RDPAP (802.31802.31802.31~{}802.31kBq/mL) was closest to the ground truth (843.36843.36843.36~{}843.36kBq/mL), suggesting that algorithms incorporating anatomical information may yield the most accurate kidney activity concentration. For both the simulated and clinical data, the order of algorithms yielding the lowest to highest noise were BSREM-RDP, BSREM-RDPAP, OSEM, and KEM; this is unsurprising as the RDP prior is designed to reduce noise. While BSREM-RDPAP has a lower noise than OSEM in each case, it also had higher uptake in the kidney: BSREM-RDPAP thus has simultaneous improvements in both bias and noise reduction. For the bone lesion, algorithms incorporating anatomical information (BSREM-RDPAP and KEM) were able to contain activity mainly inside the bone; this is demonstrated in the axial slices shown in Figure 2. These algorithms also yielded higher bone lesion activity concentration predictions than OSEM for both the simulated/clinical data, with the simulated data concentrations (347.77347.77347.77~{}347.77kBq/mL for BSREM-RDPAP and 220.56220.56220.56~{}220.56kBq/mL for KEM) closer to the ground truth (819.79819.79819.79~{}819.79kBq/mL) than OSEM (205.21205.21205.21~{}205.21kBq/mL). For the liver lesion in the simulated data, however, use of BSREM-RDPAP actually yielded a smaller activity concentration prediction (261.32261.32261.32~{}261.32kBq/mL) compared to OSEM (267.47267.47267.47~{}267.47kBq/mL). This occured because there was no anatomical difference in the surrounding tissue in the liver metastasis, while there was for the bone metastasis. These results, which are consistent with previous studies [6], suggest that anatomical priors may yield more accurate activity concentrations for bone lesion dosimetry. While this example focused primarily on organ and lesion dosimetry, PyTomography users will be to assess the merits of various reconstruction algorithms on any task specific metric.
@@
While this paper has demonstrated implementations for parallel collimator SPECT and LM-TOF PET, there exist many other modalities not presently included in the software, including (but not limited to) diverging/converging/pinhole SPECT, various forms of computed tomography, (CT), magnetic resonance (MR) imaging, and compton camera (CC) imaging. It is hoped that the modular structure and flexibility of PyTomography will encourage contributions from various research groups interested in particular imaging systems or reconstruction algorithms. Since the library is open source, newly developed functionality can be easily tested and verified by many independent research groups with their own private data. The immediate goals for future development in PyTomography at the time of publication include:
@@
    Expanded support for clinical PET data.
@@
    Support for Monte Carlo scatter correction in SPECT reconstruction.